The system of equations 4x+6y+8z=0 7x+8y+9z=0 and 3x+2y+z=0 has a) no solution b)only one solution c)two solutions d) infinite solutions
I had tried it and my ans is coming as infinite solution as rank of A matrix(4, 6,8 7,8,9 3,2,1) is 2 which is less than 3(n*n matrix holds n=3) but answer in the book is given as unique solution.

Answer by stanbon(75887)

The system of equations 4x+6y+8z=0 7x+8y+9z=0 and 3x+2y+z=0 has a) no solution b)only one solution c)two solutions d) infinite solutions
I had tried it and my ans is coming as infinite solution as rank of A matrix(4, 6,8 7,8,9 3,2,1) is 2 which is less than 3(n*n matrix holds n=3) but answer in the book is given as unique solution.
-------
Unique solution: (0,0,0)
====================
Cheers,
Stan H.